OpenStudy (anonymous):

1/5x^2 + 6/5x -8 =0 Need help factoring please!

OpenStudy (amoodarya):

factor 1/5 1/5(x^2+6x-8*5) =1/5(x^2+6x-40) =1/5(x+10)(x-4)

OpenStudy (anonymous):

so it ends up 1/5(x+10)=0 and (x-4)=0 ???

OpenStudy (anonymous):